You titrate 0.500 g of an unknown monoprotic acid with 0.0990 M NaOH. You discover that it takes 23.56 ml of base to completely neutralize the acid. what is the formula weight of the unknown acid?

If the unknown were a triprotic acid, what would its formula weight be?
Answer #1

moles of NaOH reacted = 0.099 x 23.56 / 1000 = 0.00233

if acid is monoprotic

moles of acid = 0.00233

molar mass = mass / moles = 0.500 / 0.00233

molar mass = 214.6 g/mol

if acid is triprotic acid

moles of acid reacted = 0.00233 / 3 = 0.00078 moles

molar mass = mass / moles = 0.500 / 0.00078

molar mass = 641 g/mol